# Stochastic Polyak Step-sizes and Momentum: Convergence Guarantees and Practical Performance

### v1

@@ Appendix A Further Related Work
More recently, a new line of work for adaptive algorithms has appeared, inspired by Polyak step-sizes. Some attempts for efficiently generalizing the Polyak step-size from the deterministic setting to the stochastic were made in [Rolinek and Martius, 2018, Prazeres and Oberman, 2021, Berrada et al., 2020]. Loizou et al. [2021] was the first work proposing SPSmax and providing strong convergence guarantees in different settings, including strongly convex, convex, and non-convex functions. Further extensions with decreasing variants were proposed in [Orvieto et al., 2022, Jiang and Stich, 2023]. Many recent works propose SPS-type update rules for solving optimization problems in different settings. For example, extensions of SPS to proximal setting analyzed in Schaipp et al. [2023a], SPS variants for mirror descent presented in D’Orazio et al. [2023], and SPS for federated learning proposed in Mukherjee et al. [2023]. There are also strong connections between SPS and Model-Based approaches [Asi and Duchi, 2019a, b, Chadha et al., 2022]. For further results related to SPS, see Gower et al. [2022], Li et al. [2023], Garrigos et al. [2023] and Abdukhakimov et al. [2023].
@@
In this paragraph, we review the bibliography for adaptive methods with momentum. Momentum can be either SHB or Nesterov’s acceleration, and in this case, the adaptivity can be either the step size or the momentum coefficient. In the deterministic setting, Barré et al. [2020] proposed the Accelerated Gradient Method (AGM), an adaptive algorithm built upon AGD that guarantees acceleration over GD. AGM approximates the strong convexity constant $\mu$ by the inverse of the classical Polyak Step-size. However, their algorithm still requires the knowledge of $L$ (but not $\mu$). Furthermore, Saab Jr et al. [2022] suggests an Adaptive Heavy Ball (AHB) algorithm where it approximates the constants $\mu$ and $L$ iteratively in each step based on the formula of the optimal constants of HB, but it does not show any acceleration over HB or GD. In the stochastic setting, Wang et al. [2023] propose an adaptive algorithm for a variant of SHB, named ALR-SMAG. It shows that under interpolation, ALR-SMAG enjoys a linear convergence rate for semi-strongly convex and smooth functions. Schaipp et al. [2023b] proposes a new adaptive learning rate that can be combined with any momentum-based method. Finally, Zeng et al. [2023] proposes adaptive variants for SHB in linear systems solvers.
@@
For proving convergence of SHB, Liu et al. [2020] makes the strong assumption of bounded variance and assumes knowledge of the smoothness $L$ for tuning the parameters. Another work on theoretical guarantees for SHB is Sebbouh et al. [2021], where a last iterate convergence is proved for a specific choice of $\gamma_{t}$ and $\beta_{t}$, but also requires knowledge of $L$ and only achieves rate $O(\log T/\sqrt{T})$ for the decreasing step-size variant. Furthermore, the authors of Wang et al. [2022] provide guarantees for a Polyak-based adaptive SHB method for the deterministic setting and the stochastic interpolated regime for convex and strongly convex objectives. Finally, Schaipp et al. [2023b] also assumes a stochastic convex and interpolated regime for their results.
@@ Appendix B Pseudo-codes
2: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$
3: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$ for $t=0,1,2,\dots$ do
4:      Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
5:      $\gamma_{t}=(1-\beta)\min\left\{\frac{f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}}{c\|\nabla f_{S_{t}}(x^{t})\|^{2}},\gamma_{b}\right\}$
6:      $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$
7: $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$ end for
@@
For the decreasing step-sizes, Theorems 3.5 and 3.6 make no assumption on $\beta\in[0,1)$. For SHB with MomDecSPS, the user needs to provide the momentum coefficient $\beta$, the upper bound $\gamma_{b}$, the constant $c$ and the lower bounds $\ell_{S_{t}}$, like the previous algorithm. In practice, the usual choices are $c=1$ and $\ell_{S_{t}}^{*}=0$. Note that when $t=0$, the step-size is equal to
@@
2: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$
3: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$ for $t=0,1,2,\dots$ do
4:      Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
5:      $\gamma_{t}=\min\left\{\frac{(1-\beta)[f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}]}{c_{t}\|\nabla f_{S_{t}}(x^{t})\|^{2}},\frac{\gamma_{t-1}c_{t-1}}{c_{t}}\right\}$
6:      $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$
7: $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$ end for
@@
2: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$
3: Initialization: $x^{0}\in\mathbb{R}^{d}$, $x^{-1}=x^{0}$ for $t=0,1,2,\dots$ do
4:      Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
5:      $\gamma_{t}=\min\left\{\frac{(1-\beta)[f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}]}{c\|\nabla f_{S_{t}}(x^{t})\|^{2}\sqrt{\sum_{s=0}^{t}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}},\gamma_{t-1}\right\}$
6:      $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$
7: $x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$ end for
@@ C.1 Basic Definitions
$f(x)-f(y)\leq\langle\nabla f(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}$
@@ C.3 Iterate Moving Average
As mentioned in the main paper, the iterates of the SHB method are equal to the IMA method. The following proposition is from Sebbouh et al. [2021]. We include the proof for completeness.
@@
then the $x_{t}$ iterates of the IMA method are equal to the $x_{t}$ iterates produced by the SHB method.
@@
We will start from the IMA update rule and we will prove that its $(x^{t})$ iterates satisfy the SHB update rule. Suppose that we have the $(z^{t+1})$ and $(x^{t+1})$ iterates as well as the $(\lambda_{t})$ and $(\eta_{t})$ sequences. Define the sequences $(\beta_{t})$ and $(\gamma_{t})$ using eqs. 9 and 10. We will show that $(x^{t})$ iterates satisfy the SHB update rule. Solving for $z^{t+1}$ in eq. 8 we get
@@
=(7)${=}\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}\left(z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t})\right)$ )
=(11)${=}\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}\left(x^{t}+\lambda_{t}(x^{t}-x^{t-1})-\eta_{t}\nabla f_{S_{t}}(x^{t}))\right)$ )
@@
=(9,10)${=}x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta_{t}(x^{t}-x^{t-1}),$ ) ,
@@ Appendix D Proofs of the main results
In this section we present the proofs of the main theoretical results presented in the main paper, i.e., the convergence analysis of SHB with MomSPSmax, MomDecSPS and MomAdaSPS for convex and smooth functions $f_{i}$ and $f$ of Problem 1. The idea for all the proofs is that we firstly show the corresponding result in the IMA setting, and then we transfer it to the SHB setting with the help of Proposition 2.1.
Compared to the analysis of SGD with SPS, the SHB update rule requires taking into consideration extra terms related to the previous iterate $x^{t-1}$. This needs new algebraic tricks in order to use convexity, and after that, one needs to deal with the two quantities $f(x^{t})$ and $f(x^{t-1})$ at the same time. We deal with this using the IMA framework, similarly with Sebbouh et al. [2021]. However, we highlight that we finish our proofs with a different approach. More specifically, in the proof of Theorem G.1 in Sebbouh et al. [2021] the authors choose the momentum coefficients $\lambda_{t}$ in such a way that the main sum telescopes and the final result only has the expectation of the last iterate. In our proof, we make no such simplifications, and instead, we use Jensen’s inequality to finish. This approach allows us to have more freedom in the selection of $\lambda_{t}$. This is more prominent in the proofs of the decreasing variants.
@@
First, we state and prove the corresponding theorem for SPSmax in the IMA setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of IMA with
@@
[…]
@@
[…]
@@
$\overset{z^{0}=x^{0}}{\leq}\|x^{0}-x^{*}\|^{2}+\sum_{t=0}^{T-1}2\left(\eta_{b}\left(1+\lambda_{t}\right)-\alpha\left(\lambda_{t}+\frac{1}{2}\right)\right)\sigma^{2}.$ (14)
@@
$=2\alpha\left(\lambda_{T-1}+\frac{1}{2}\right)\operatorname{\mathbb{E}}[f(x^{T-1})-f(x^{*})]+\sum_{t=0}^{T-2}2\alpha\left(\lambda_{t}+\frac{1}{2}\right)\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-\sum_{t=0}^{T-2}2\eta_{b}\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\geq\left[2\alpha\left(\lambda_{T-1}+\frac{1}{2}\right)-2\eta_{b}\lambda_{T}\right]\operatorname{\mathbb{E}}[f(x^{T-1})-f(x^{*})]+\sum_{t=0}^{T-2}\left[2\alpha\left(\lambda_{t}+\frac{1}{2}\right)-2\eta_{b}\lambda_{t+1}\right]\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\sum_{t=0}^{T-1}\left[2\alpha\left(\lambda_{t}+\frac{1}{2}\right)-2\eta_{b}\lambda_{t+1}\right]\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]\leq\|x^{0}-x^{*}\|^{2}+\sum_{t=0}^{T-1}2\left(\eta_{b}\left(1+\lambda_{t}\right)-\alpha\left(\lambda_{t}+\frac{1}{2}\right)\right)\sigma^{2}.$ (16)
In eq. 16, setting $\lambda_{t}=\lambda,\forall t>0$ with $0\leq\lambda<\frac{\alpha}{2(\eta_{b}-\alpha)}$ (in order to ensure that quantity $2\alpha\left(\lambda+\frac{1}{2}\right)-2\eta_{b}\lambda$ in the LHS is positive) we get
@@
Here we state and prove the corresponding theorem for DecSPS in the IMA setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of IMA with
@@
$2\sum_{t=0}^{T-1}\langle\nabla f_{S_{t}}(x^{t}),z^{t}-x^{*}\rangle$ $\leq\sum_{t=0}^{T-1}\left[\frac{\|z^{t}-x^{*}\|^{2}}{\eta_{t}}-\frac{\|z^{t+1}-x^{*}\|^{2}}{\eta_{t}}+\eta_{t}\|\nabla f_{S_{t}}(x^{t})\|^{2}\right]$
@@
$2\sum_{t=0}^{T-1}\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{t}),z^{t}-x^{*}\rangle\right]$ =$\overset{(z^{0}=x^{0})}{=}2\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{0}),x^{0}-x^{*}\rangle\right]+2\sum_{t=1}^{T-1}\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{t}),z^{t}-x^{*}\rangle\right]$
[…]
$\overset{convexity}{\geq}2\operatorname{\mathbb{E}}\left[f(x^{0})-f(x^{*})\right]+\sum_{t=1}^{T-1}\operatorname{\mathbb{E}}\left[2(f(x^{t})-f(x^{*}))+2\lambda_{t}(f(x^{t})-f(x^{t-1}))\right]$
@@
$\quad-\sum_{t=1}^{T-2}2\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\quad+\sum_{t=1}^{T-2}2\lambda_{t}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-\sum_{t=1}^{T-2}2\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
Here we state and prove the corresponding theorem for AdaSPS in the IMA setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of IMA with
@@
$2\sum_{t=0}^{T-1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-2\lambda_{1}[f(x^{0})-f(x^{*})]$ $\leq\operatorname{\mathbb{E}}\left[cLD^{2}_{z}\sqrt{\sum_{s=0}^{T-1}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}+\frac{1}{c}\sqrt{\sum_{s=0}^{T-1}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}\right]$

### v2

@@ Appendix A Further Related Work
More recently, a new line of work for adaptive algorithms has appeared, inspired by Polyak step-sizes. Some attempts for efficiently generalizing the Polyak step-size from the deterministic setting to the stochastic were made in Rolinek and Martius [2018], Prazeres and Oberman [2021], Berrada et al. [2020]. Loizou et al. [2021] was the first work proposing SPSmax and providing strong convergence guarantees in different settings, including strongly convex, convex, and non-convex functions. Further extensions with decreasing variants were proposed in Orvieto et al. [2022], Jiang and Stich [2023]. Many recent works propose SPS-type update rules for solving optimization problems in different settings. For example, extensions of SPS to proximal setting analyzed in Schaipp et al. [2023a], SPS variants for mirror descent presented in D’Orazio et al. [2023], and SPS for federated learning proposed in Mukherjee et al. [2023]. There are also strong connections between SPS and Model-Based approaches [Asi and Duchi, 2019a, b, Chadha et al., 2022]. For further results related to SPS, see Gower et al. [2022], Li et al. [2023], Garrigos et al. [2023] and Abdukhakimov et al. [2023].
@@
In this paragraph, we review the bibliography for adaptive methods with momentum. Momentum can be either SHB or Nesterov’s acceleration, and in this case, the adaptivity can be either the step-size or the momentum coefficient. In the deterministic setting, Barré et al. [2020] proposed the Accelerated Gradient Method (AGM), an adaptive algorithm built upon AGD that guarantees acceleration over GD. AGM approximates the strong convexity constant $\mu$ by the inverse of the classical Polyak Step-size. However, their algorithm still requires the knowledge of $L$ (but not $\mu$). Furthermore, Saab Jr et al. [2022] suggests an Adaptive Heavy Ball (AHB) algorithm where it approximates the constants $\mu$ and $L$ iteratively in each step based on the formula of the optimal constants of HB, but it does not show any acceleration over HB or GD. In the stochastic setting, Wang et al. [2023] propose an adaptive algorithm for a moving averaged gradient momentum, named ALR-SMAG. It shows that ALR-SMAG enjoys a linear convergence rate for semi-strongly convex and smooth functions. Schaipp et al. [2023b] proposes a new adaptive learning rate that can be combined with any momentum-based method. Finally, Zeng et al. [2023] proposes adaptive variants for SHB in linear systems solvers.
@@
For proving convergence of SHB, Liu et al. [2020] makes the strong assumption of bounded variance and assumes knowledge of the smoothness $L$ for tuning the parameters. Another work on theoretical guarantees for SHB is Sebbouh et al. [2021], where a last iterate convergence is proved for a specific choice of $\gamma_{t}$ and $\beta_{t}$, but also requires knowledge of $L$ and only achieves rate $O(\log T/\sqrt{T})$ for the decreasing step-size variant. Furthermore, the authors of Wang et al. [2022] propose a Polyak-based stepsize for a moving averaged gradient momentum variant, different from SHB, and provide guarantees for the deterministic and stochastic setting for strongly convex objectives. Finally, Schaipp et al. [2023b] assumes a stochastic convex and interpolated regime for their results.
@@ Appendix B Pseudo-codes
[…]
3:Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
4:$\gamma_{t}=(1-\beta)\min\left\{\frac{f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}}{c\|\nabla f_{S_{t}}(x^{t})\|^{2}},\gamma_{b}\right\}$
5:$x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$EndFor
@@
For the decreasing step-sizes, Theorems 3.6 and 3.7 make no assumption on $\beta\in[0,1)$. For SHB with MomDecSPS, the user needs to provide the momentum coefficient $\beta$, the upper bound $\gamma_{b}$, the constant $c$ and the lower bounds $\ell_{S_{t}}$, like the previous algorithm. In practice, the usual choices are $c=1$ and $\ell_{S_{t}}^{*}=0$. Note that when $t=0$, the step-size is equal to
@@
[…]
3:Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
4:$\gamma_{t}=\min\left\{\frac{(1-\beta)[f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}]}{c_{t}\|\nabla f_{S_{t}}(x^{t})\|^{2}},\frac{\gamma_{t-1}c_{t-1}}{c_{t}}\right\}$
5:$x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$EndFor
@@
[…]
3:Choose uniformly at random $S_{t}\subseteq\{1,\dots,n\}$
4:$\gamma_{t}=\min\left\{\frac{(1-\beta)[f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}]}{c\|\nabla f_{S_{t}}(x^{t})\|^{2}\sqrt{\sum_{s=0}^{t}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}},\gamma_{t-1}\right\}$
5:$x^{t+1}=x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta(x^{t}-x^{t-1})$EndFor
@@ C.1 Basic Definitions
$|f(x)-f(y)-\langle\nabla f(y),x-y\rangle|\leq\frac{L}{2}\|x-y\|^{2}$
@@ C.3 Iterate Moving Average
As mentioned in the main paper, the iterates of the SHB method are equal to the 2.3 method. The following proposition is from Sebbouh et al. [2021]. We include the proof for completeness.
@@
then the $x_{t}$ iterates of the 2.3 method are equal to the $x_{t}$ iterates produced by the SHB method.
@@
We will start from the 2.3 update rule and we will prove that its $(x^{t})$ iterates satisfy the SHB update rule. Suppose that we have the $(z^{t+1})$ and $(x^{t+1})$ iterates as well as the $(\lambda_{t})$ and $(\eta_{t})$ sequences. Define the sequences $(\beta_{t})$ and $(\gamma_{t})$ using eqs. 9 and 10. We will show that $(x^{t})$ iterates satisfy the SHB update rule. Solving for $z^{t+1}$ in eq. 8 we get
@@
${=}\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}\left(z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t})\right)$
${=}\frac{\lambda_{t+1}}{\lambda_{t+1}+1}x^{t}+\frac{1}{\lambda_{t+1}+1}\left(x^{t}+\lambda_{t}(x^{t}-x^{t-1})-\eta_{t}\nabla f_{S_{t}}(x^{t}))\right)$
@@
${=}x^{t}-\gamma_{t}\nabla f_{S_{t}}(x^{t})+\beta_{t}(x^{t}-x^{t-1}),$
@@ Appendix D Proofs of the main results
In this section we present the proofs of the main theoretical results presented in the main paper, i.e., the convergence analysis of SHB with MomSPSmax, MomDecSPS and MomAdaSPS for convex and smooth functions $f_{i}$ and $f$ of Problem 1. The idea for all the proofs is that we firstly show the corresponding result in the 2.3 setting, and then we transfer it to the SHB setting with the help of Proposition 2.1.
Compared to the analysis of SGD with SPS, the SHB update rule requires taking into consideration extra terms related to the previous iterate $x^{t-1}$. This needs new algebraic tricks in order to use convexity, and after that, one needs to deal with the two quantities $f(x^{t})$ and $f(x^{t-1})$ at the same time. We deal with this using the 2.3 framework, similarly with Sebbouh et al. [2021]. However, we highlight that we finish our proofs with a different approach. More specifically, in the proof of Theorem G.1 in Sebbouh et al. [2021] the authors choose the momentum coefficients $\lambda_{t}$ in such a way that the main sum telescopes and the final result only has the expectation of the last iterate. In our proof, we make no such simplifications, and instead, we use Jensen’s inequality to finish. This approach allows us to have more freedom in the selection of $\lambda_{t}$. This is more prominent in the proofs of the decreasing variants.
@@
First, we state and prove the corresponding theorem for SPSmax in the 2.3 setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of 2.3 with
@@
${=}\|z^{t}-x^{*}\|^{2}-2\eta_{t}\langle\nabla f_{S_{t}}(x^{t}),x^{t}-x^{*}\rangle-2\eta_{t}\lambda_{t}\langle\nabla f_{S_{t}}(x^{t}),x^{t}-x^{t-1}\rangle$
@@
${\leq}\|z^{t}-x^{*}\|^{2}-2\alpha\left(\lambda_{t}+\frac{1}{2}\right)[f_{S_{t}}(x^{t})-\ell_{S_{t}}^{*}]+2\eta_{b}\lambda_{t}[f_{S_{t}}(x^{t-1})-\ell_{S_{t}}^{*}]$
@@
${\leq}\|x^{0}-x^{*}\|^{2}+\sum_{t=0}^{T-1}2\left(\eta_{b}\left(1+\lambda_{t}\right)-\alpha\left(\lambda_{t}+\frac{1}{2}\right)\right)\sigma^{2}.$ (14)
@@
$=2\alpha\left(\lambda_{T-1}+\frac{1}{2}\right)\operatorname{\mathbb{E}}[f(x^{T-1})-f(x^{*})]+\sum_{t=0}^{T-2}2\alpha\left(\lambda_{t}+\frac{1}{2}\right)\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
$\quad-\sum_{t=0}^{T-2}2\eta_{b}\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\geq\left[2\alpha\left(\lambda_{T-1}+\frac{1}{2}\right)-2\eta_{b}\lambda_{T}\right]\operatorname{\mathbb{E}}[f(x^{T-1})-f(x^{*})]$
$\quad+\sum_{t=0}^{T-2}\left[2\alpha\left(\lambda_{t}+\frac{1}{2}\right)-2\eta_{b}\lambda_{t+1}\right]\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\sum_{t=0}^{T-1}\left[2\alpha\left(\lambda_{t}+\frac{1}{2}\right)-2\eta_{b}\lambda_{t+1}\right]\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
$\leq\|x^{0}-x^{*}\|^{2}+\sum_{t=0}^{T-1}2\left(\eta_{b}\left(1+\lambda_{t}\right)-\alpha\left(\lambda_{t}+\frac{1}{2}\right)\right)\sigma^{2}.$ (16)
In section D.1, setting $\lambda_{t}=\lambda,\forall t>0$ with $0\leq\lambda<\frac{\alpha}{2(\eta_{b}-\alpha)}$ (in order to ensure that quantity $2\alpha\left(\lambda+\frac{1}{2}\right)-2\eta_{b}\lambda$ in the LHS is positive) we get
@@
Here we state and prove the corresponding theorem for DecSPS in the 2.3 setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of 2.3 with
@@
$2\sum_{t=0}^{T-1}\langle\nabla f_{S_{t}}(x^{t}),z^{t}-x^{*}\rangle$
$\leq\sum_{t=0}^{T-1}\left[\frac{\|z^{t}-x^{*}\|^{2}}{\eta_{t}}-\frac{\|z^{t+1}-x^{*}\|^{2}}{\eta_{t}}+\eta_{t}\|\nabla f_{S_{t}}(x^{t})\|^{2}\right]$
@@
$2\sum_{t=0}^{T-1}$ $\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{t}),z^{t}-x^{*}\rangle\right]$
${=}2\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{0}),x^{0}-x^{*}\rangle\right]+2\sum_{t=1}^{T-1}\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{t}),z^{t}-x^{*}\rangle\right]$
${=}2\operatorname{\mathbb{E}}\left[\langle\nabla f(x^{0}),x^{0}-x^{*}\rangle\right]+\sum_{t=1}^{T-1}\operatorname{\mathbb{E}}\left[2\langle\nabla f(x^{t}),x^{t}-x^{*}\rangle+2\lambda_{t}\langle\nabla f(x^{t}),x^{t}-x^{t-1}\rangle\right]$
${\geq}2\operatorname{\mathbb{E}}\left[f(x^{0})-f(x^{*})\right]+\sum_{t=1}^{T-1}\operatorname{\mathbb{E}}\left[2(f(x^{t})-f(x^{*}))+2\lambda_{t}(f(x^{t})-f(x^{t-1}))\right]$
@@
$\quad-\sum_{t=0}^{T-2}2\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]$
@@
$\quad+\sum_{t=1}^{T-2}2\lambda_{t}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-\sum_{t=1}^{T-2}2\lambda_{t+1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-2\lambda_{1}\operatorname{\mathbb{E}}[f(x^{0})-f(x^{*})]$
@@
Here we state and prove the corresponding theorem for AdaSPS in the 2.3 setting.
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of 2.3 with
@@
$2\sum_{t=0}^{T-1}\operatorname{\mathbb{E}}[f(x^{t})-f(x^{*})]-2\lambda_{1}[f(x^{0})-f(x^{*})]$
$\leq\operatorname{\mathbb{E}}\left[cLD^{2}_{z}\sqrt{\sum_{s=0}^{T-1}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}+\frac{1}{c}\sqrt{\sum_{s=0}^{T-1}f_{S_{s}}(x^{s})-\ell_{S_{s}}^{*}}\right]$